OpenStudy (anonymous):

simplify (2)(1 + (1i)/(n))

OpenStudy (primeralph):

write the question better

OpenStudy (anonymous):

okay hold on \[2 \times 1 + \frac{ 1i }{ n}\]

OpenStudy (primeralph):

multiply 2 by each one insde

OpenStudy (anonymous):

i get 2 + 2i/n is that correct?

OpenStudy (anonymous):

@primeralph are u there?

OpenStudy (primeralph):

yeah

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

@primeralph are u there?

OpenStudy (anonymous):

i get 2 + 2i/n is that correct?

OpenStudy (anonymous):

okay hold on \[2 \times 1 + \frac{ 1i }{ n}\]

OpenStudy (primeralph):

correct if there are parentheses